# New phenomena arising from ${\mathcal{L}}$-invariants of modular forms

### v1

@@ 1.1. Archimedean distributions of Hecke eigenvalues
for a fixed eigenform $f$ and $p$ ranging over the primes. Its importance is based on connections, independently made by Sato and Tate, to distributions of point-counts for the reductions of rational elliptic curves over finite fields. (See [71, p. 106-107].) Its proof is also a landmark in modern number theory, for it was driven by potential modularity techniques invented by Taylor and his collaborators, generalizing the modularity theorems of Wiles, Taylor and Wiles, and of Breuil, Conrad, Diamond, and Taylor [76, 72, 16].
@@ 1.2. Non-Archimedean distributions of Hecke eigenvalues
$\mathbf{x}_{T}=\left\{\frac{p+1}{k}\cdot v_{p}(a_{p}(f))~{}\Big{|}~{}\text{$f$is an $\Gamma_{0}$-eigenform of weight $k\leq T$}\right\}.$
@@
Figure 1. $2$-adic slopes in level $\operatorname{SL}_{2}({\mathbb{Z}})$ for $k\leq 512$. (Total data: 5334 eigenforms.)
@@
Figure 2. Left: Normalized $2$-adic slopes in level $\operatorname{SL}_{2}({\mathbb{Z}})$ for $k\leq 512$.
@@
with $52\leq a\leq p-$ 5 and $b\in{\mathbb{Z}}$. Then, $\mathbf{x}_{T}(\overline{\rho})$ becomes equidistributed on $[0,1]$ for Lebesgue measure as $T\rightarrow\infty$.
@@
The proof of Theorem 1.2 is based on a proof of a more remarkable theorem on regular mod $p$ Galois representations. Namely, we developed, between 2015 and 2019, a combinatorial model for slopes of $a_{p}$ that is called the ghost series. The ghost conjecture asserts that the ghost series predicts the slopes of weight $k$ eigenforms, via its Newton polygons, whenever $\overline{\rho}$ is regular [10, 11]. The main theorem of Liu–Truong–Xiao–Zhao is that the ghost conjecture holds under the hypotheses listed in Theorem 1.2. See [52, Theorem 1.3]. In this way, analyses of slopes are reduced to analyses for the combinatorially-defined ghost series, which are easier. In fact, the authors themselves proved Gouvêa’s distribution conjecture is a corollary of their ghost conjecture. See [11, Corollary 3.3]. Nonetheless, all but the smallest fraction of credit for Theorem 1.2 should be directed at Liu, Truong, Xiao, and Zhao, for their amazing work proving the ghost conjecture.
@@
One might also seek a conceptual understanding of the normalizing factor $\frac{p+1}{k}$, since we have such an understanding of the factor $p^{\frac{1-k}{2}}$ in the Archimedean cases. Normalizing by $k$ seems reasonable enough because there is a linear-in-$k$ number of eigenforms of weight $k$. Dividing by $k$ is also natural if one works at level $\Gamma_{0}(pM)$, where the largest $a_{p}$-slope is no more than $k-1$. (Gouvêa technically normalizes by $\frac{p+1}{k-1}$, even.) So, it is the factor of $p+1$ that remains obscure. Why is $v_{p}(a_{p}(f))\leq\frac{k}{p+1}$ almost all the time? We emphasize “almost” here, as the proposed statistical bound does not always hold. After all, it can happen that $a_{p}(f)=0$. (Calegari and Sadrari did recently prove that for fixed $M$ there are a finite number of eigenforms with vanishing $a_{p}$, excluding the forms with complex multiplication. See [23, Theorem 1.1].)
@@ 1.3. The ${\mathcal{L}}$-invariants
    If $p$ divides $M$ exactly once and $f$ has weight $k$, then $a_{p}(f)=\pm p^{\frac{k}{2}-1}$. Even more, the number of occurrences of $a_{p}(f)>0$ versus $a_{p}(f)<0$ in any given weight $k$ is bounded. (For this bounded-ness, see [53, Section 3] when $M$ is squarefree, as well as Theorem 7.4.)
@@
These ${\mathcal{L}}$-invariants are quickly calculated. For instance, if $j=j_{E}$ is the $j$-invariant of $E$, then $j^{-1}\in p{\mathbb{Z}}_{p}$ and $q_{E}$ can be expressed as an infinite series (see [56, II.1])
@@
be the semi-simple (irreducible, even) $p$-adic Galois representation attached to $f$. Let $r_{f}:=\rho_{f}|_{\operatorname{Gal}(\overline{{{\mathbb{Q}}}}_{p}/{\mathbb{Q}}_{p})}$ be its restriction to a decomposition group at $p$. Then:
@@ 1.4. Prior computations
Despite the apparent interest in ${\mathcal{L}}$-invariants, with many researchers proposing different definitions and perspectives, there has been a definite deficit on explicitly computing them. Exactly two examples not coming from elliptic curves were computed in the seminal work of Mazur, Tate, and Teitelbaum. See [56, pg. 46]. [^1] The second example appears to be off by a factor of $p=5$! (Exclamation, not factorial.) In the decades following that study, only a few more examples were computed, in various works of Coleman and Teitelbaum, of Coleman, Stevens, and Teitelbaum, and of Lauder [27, 26, 50].
@@
The works of Gräf and Anni–Böckle–Gräf–Troya have limitations. For instance, they “only” calculate the list of ${\mathcal{L}}$-invariants arising in a fixed weight $k$ and with a fixed sign $a_{p}(f)=\pm p^{\frac{k}{2}-1}$. They cannot say which ${\mathcal{L}}$-invariant corresponds to which form. Their works also have distinct, beneficial, features. Gräf’s work computes a little more than ${\mathcal{L}}$-invariants, since it also calculates a natural linear operator for which the ${\mathcal{L}}$-invariants are eigenvalues. (This is a special feature of Teitelbaum’s definition.) The method via the Greenberg–Stevens formula notably presents the chance to perform calculations at level prime-to-$p$ as well. The ratio $-2a_{p}^{\prime}(k)/a_{p}(f)$ is still called an ${\mathcal{L}}$-invariant then, however it is a purely global (even in Galois-theoretic terms) invariant. See [47, 57], for instance. Such calculations could be relevant to questions on constant slope families raised in [6]. To our knowledge, no one has systematically pursued such calculations.
@@ 1.5. The exceptional zero method
One may overcome this by replacing $f$ by a twist $f_{\chi}$ by a quadratic Dirchlet character $\chi$ trivial at $p$. The twist doesn’t affect the ${\mathcal{L}}$-invariant, and $f_{\chi}$ will have a non-vanishing central $L$-value for some $\chi$. Incidentally, twisting also allows us to use (EZ) when $a_{p}$ is negative: if $a_{p}(f)<0$, we could perform a twist by a character where $\chi(p)=-1$, instead. See Section 2.3 for a precise discussion.
@@ 1.6. Data analysis
Figure 3. $2$-adic valuations of ${\mathcal{L}}$-invariants in level $\Gamma_{0}(2)$ with respect to weights $10\leq k\leq 804$. (Total data: 13465 eigenforms.)
@@
modular of level $\Gamma_{0}(N)$, and its local restriction $\overline{r}=\overline{\rho}|_{\mathrm{Gal}(\overline{{{\mathbb{Q}}}}_{p}/{\mathbb{Q}}_{p})}$. For simplicity, assume $\overline{r}$ has only scalar endomorphisms. (See Section 6.2 for further discussion.) Given $\overline{r}$, a weight $k\geq 2$, and a sign $\pm$, we consider the set $X_{k}^{\pm}(\overline{r})\subseteq{\mathbb{P}}^{1}(\overline{{{\mathbb{Q}}}}_{p})$ consisting of ${\mathcal{L}}$-invariants that arise from semi-stable, but non-crystalline, lifts of $\overline{r}$ with weight $k$ and sign $\pm$. The set $X_{k}^{\pm}(\overline{r})$ is the ¯$\overline{{{\mathbb{Q}}}}_{p}$-points of the rigid generic fiber of a certain deformation ring, embedded into ${\mathbb{P}}^{1}$ by means of the ${\mathcal{L}}$-invariant.
@@ 1.7. The distributional conjecture
[…]
Then, $\mathbf{y}_{T}^{\pm}(\overline{\rho})$ becomes equidistributed for Lebesgue measure on $[-1,0]$ as $T\rightarrow\infty$.
When $p=2$, the normalization in Conjecture 1.3 is by $\frac{6}{k}$. (The slope in Figure 3 is about $-\frac{1}{6}$.) Figure 4 presents a scatter plot and histogram to support Conjecture 1.3. In level $\operatorname{SL}_{2}({\mathbb{Z}})$, the only mod $p=2$ Galois representation is $\overline{\rho}=1\oplus 1$. We do not filter by the sign $\pm$ in the plot itself. See Section 8.3 for more plots.
@@ 1.10. Acknowledgments
The research reported on here was carried out, partially, while the authors visited the Max-Planck Institute for Mathematics in Bonn, Germany. The staff and members of the institute are warmly thanked for their support and hospitality. Bergdall’s research has been partially supported by Simons Foundation Travel Support Grant for Mathematicians #713782 and National Science Foundation grant DMS-2302284. Pollack’s research has been partially supported by National Science Foundation grants DMS-1702178 and DMS-2302285 and by Simons Foundation Travel Support Grant for Mathematicians MPS-TSM-00002405.
@@ 2.1. Notations
We fix an isomorphism $\iota:{\mathbb{C}}\simeq\overline{{{\mathbb{Q}}}}_{p}$, thus allowing us to view an eigenform as defined over ¯$\overline{{{\mathbb{Q}}}}_{p}$. On ¯$\overline{{{\mathbb{Q}}}}_{p}$, we also fix a $p$-adic valuation $v_{p}$ such that $v_{p}(p)=1$. We always measure $p$-adic numbers with $v_{p}(-)$. If we apply $v_{p}(-)$ to a complex number, we mean to implicitly apply $\iota$ first.
@@ 2.2. Modular symbols and $p$-adic $L$-functions
[…]
where $r\in{\mathbb{Q}}$ and $P(z)$ is an polynomial over ${\mathbb{Z}}$ of degree at most $k-2$ (as in [56]). The modular integrals of an eigenform $f$ may be decomposed into plus and minus parts and renormalized to take algebraic values:
@@
[…]
for all $P(z)$ of degree less than or equal to $k-2$. The distribution $\mu_{f}$ is the $p$-adic $L$-function of $f$.
We note that $p$-adic $L$-functions can also be defined for eigenforms that are old at $p$, but in that case one also needs to choose a root of the Hecke polynomial $x^{2}-a_{p}(f)x+p^{k-1}$. When this root has valuation strictly less than $k-1$, then the construction of $\mu_{f}$ follows the same approach except that the definition of $\mu_{f}$ in (2) is slightly different and depends upon this choice of root. We will not discuss $p$-adic $L$-functions of forms that are old at $p$ any further.
Let us describe the $p$-adic $L$-function in the “$s$-variable”. First, write each $x\in{\mathbb{Z}}_{p}$ uniquely as $x=\langle x\rangle\omega(x)$, where $\langle x\rangle\in 1+p{\mathbb{Z}}_{p}$ and $\omega(x)$ is a $(p-1)$-st root of unity. Then, for a tame character $\psi:{\mathbb{Z}}_{p}^{\times}\to{\mathbb{F}}_{p}^{\times}\to\overline{{{\mathbb{Q}}}}_{p}^{\times}$, we set
[…]
This is the “$\psi$-branch” of the $p$-adic $L$-function and here $\psi$ has $p-1$ choices: $1$, $\omega$, $\omega^{2}$, $\dots$, $\omega^{p-2}$. Nonetheless, for $j$ an integer, we often abuse notation and simply write
[…]
@@
[…]
@@
[…]
@@
[…]
@@
This divisibility is almost [56, pg. 13, III], but in that reference a lower bound of $n\left(m-\frac{k-2}{2}\right)-\frac{k-2}{2}$ is given. However, in our situation of a form whose level is divisible by $p$ exactly once, we can achieve the slightly stronger bound of this lemma. Indeed, if $f$ had level prime-to-$p$ and $\alpha$ was a root of the Hecke polynomial of $f$ at $p$, then the very definition of $\mu_{f}$ yields that $\alpha^{n+1}\mu_{f}(P,a,p^{n})$ is integral. However, in our case, (2) implies that $\alpha^{n}\mu_{f}(P,a,p^{n})$ is integral where $\alpha=a_{p}(f)=\pm p^{\frac{k}{2}-1}$. Tracing through the proof of [56, pg. 13, III] using this stronger fact yields the claim of this lemma. ∎

### v2

@@ 1.1. Archimedean distributions of Hecke eigenvalues
for a fixed eigenform $f$ and $p$ ranging over the primes. Its importance is based on connections, independently made by Sato and Tate, to distributions of point-counts for the reductions of rational elliptic curves over finite fields. (See [72, p. 106-107].) Its proof is also a landmark in modern number theory, for it was driven by potential modularity techniques invented by Taylor and his collaborators, generalizing the modularity theorems of Wiles, Taylor and Wiles, and of Breuil, Conrad, Diamond, and Taylor [77, 73, 16].
@@ 1.2. Non-Archimedean distributions of Hecke eigenvalues
$\mathbf{x}_{T}=\left\{\frac{p+1}{k}\cdot v_{p}(a_{p}(f))~\Big|~\text{$f$ is a $\Gamma_{0}$-eigenform of weight $k\leq T$}\right\}.$
@@
Figure 1. 22-adic slopes in level $\operatorname{SL}_{2}({\mathbb{Z}})$ for $k\leq 512$. (Total data: 5334 eigenforms.)
@@
Figure 2. Left: Normalized 22-adic slopes in level $\operatorname{SL}_{2}({\mathbb{Z}})$ for $k\leq 512$.
@@
with $2\leq a\leq p-5$ and $b\in{\mathbb{Z}}$. Then, $\mathbf{x}_{T}(\overline{\rho})$ becomes equidistributed on $[0,1]$ for Lebesgue measure as $T\rightarrow\infty$.
@@
The proof of Theorem 1.2 is based on a proof of a more remarkable theorem on regular mod $p$ Galois representations. Namely, we developed, between 2015 and 2019, a combinatorial model for slopes of $a_{p}$ that is called the ghost series. The ghost conjecture asserts that the ghost series predicts the slopes of weight $k$ eigenforms, via its Newton polygons, whenever $\overline{\rho}$ is regular [10, 11]. The main theorem of Liu–Truong–Xiao–Zhao is that the ghost conjecture holds under the hypotheses listed in Theorem 1.2. See [51, Theorem 1.3]. In this way, analyses of slopes are reduced to analyses for the combinatorially-defined ghost series, which are easier. In fact, the authors themselves proved Gouvêa’s distribution conjecture is a corollary of their ghost conjecture. See [11, Corollary 3.3]. Nonetheless, all but the smallest fraction of credit for Theorem 1.2 should be directed at Liu, Truong, Xiao, and Zhao, for their amazing work proving the ghost conjecture.
@@
One might also seek a conceptual understanding of the normalizing factor $\frac{p+1}{k}$, since we have such an understanding of the factor $p^{\frac{1-k}{2}}$ in the Archimedean cases. Normalizing by $k$ seems reasonable enough because there is a linear-in-$k$ number of eigenforms of weight $k$. Dividing by $k$ is also natural if one works at level $\Gamma_{0}(pM)$, where the largest $a_{p}$-slope is no more than k1k-1. (Gouvêa technically normalizes by $\frac{p+1}{k-1}$, even.) So, it is the factor of $p+1$ that remains obscure. Why is $v_{p}(a_{p}(f))\leq\frac{k}{p+1}$ almost all the time? We emphasize “almost” here, as the proposed statistical bound does not always hold. After all, it can happen that $a_{p}(f)=0$. (Calegari and Sadrari did recently prove that for fixed $M$ there are a finite number of eigenforms with vanishing $a_{p}$, excluding the forms with complex multiplication. See [23, Theorem 1.1].)
@@ 1.3. The $\mathcal{L}$}-invariants
    If $p$ divides $M$ exactly once and $f$ has weight $k$, then $a_{p}(f)=\pm p^{\frac{k}{2}-1}$. Even more, the number of occurrences of $a_{p}(f)>0$ versus $a_{p}(f)<0$ in any given weight $k$ is bounded. (For this bounded-ness, see [54, Corollary 1.4] (or [53, Section 3] if $M$ is squarefree) or the Theorem 7.4 of Anni, Ghitza, and Medvedovsky.)
@@
These $\mathcal{L}$}-invariants are quickly calculated. For instance, if $j=j_{E}$ is the $j$-invariant of $E$, then $j^{-1}\in p{\mathbb{Z}}_{p}$ and $q_{E}$ can be expressed as an infinite series (see [57, II.1])
@@
be the semi-simple (irreducible in fact, since $f$ is cuspidal) $p$-adic Galois representation attached to $f$. Let $r_{f}:=\rho_{f}|_{\operatorname{Gal}(\overline{{{\mathbb{Q}}}}_{p}/{\mathbb{Q}}_{p})}$ be its restriction to a decomposition group at $p$. Then:
@@ 1.4. Prior computations
Despite the apparent interest in $\mathcal{L}$}-invariants, with many researchers proposing different definitions and perspectives, there has been a definite deficit on explicitly computing them. Exactly two examples not coming from elliptic curves were computed in the seminal work of Mazur, Tate, and Teitelbaum. See [57, pg. 46]. [^1] The second example appears to be off by a factor of p=5p=5! (Exclamation, not factorial.) In the decades following that study, only a few more examples were computed, in various works of Coleman and Teitelbaum, of Coleman, Stevens, and Teitelbaum, and of Lauder [27, 26, 50].
@@
The works of Gräf and Anni–Böckle–Gräf–Troya have limitations. For instance, they “only” calculate the list of $\mathcal{L}$}-invariants arising in a fixed weight $k$ and with a fixed sign $a_{p}(f)=\pm p^{\frac{k}{2}-1}$. They cannot say which $\mathcal{L}$}-invariant corresponds to which form. Their works also have distinct, beneficial, features. Gräf’s work computes a little more than $\mathcal{L}$}-invariants, since it also calculates a natural linear operator for which the $\mathcal{L}$}-invariants are eigenvalues. (This is a special feature of Teitelbaum’s definition.) The method via the Greenberg–Stevens formula notably presents the chance to perform calculations at level prime-to-$p$ as well. The ratio $2a_{p}^{\prime}(k)/a_{p}(f)$ is still called an $\mathcal{L}$}-invariant then, however it is a purely global (even in Galois-theoretic terms) invariant. See [47, 58], for instance. Such calculations could be relevant to questions on constant slope families raised in [6]. To our knowledge, no one has systematically pursued such calculations.
@@ 1.5. The exceptional zero method
One may overcome this by replacing $f$ by a twist $f_{\chi}$ by a quadratic Dirichlet character $\chi$ trivial at $p$. The twist doesn’t affect the $\mathcal{L}$}-invariant, and $f_{\chi}$ will have a non-vanishing central $L$-value for some $\chi$. Incidentally, twisting also allows us to use (EZ) when $a_{p}$ is negative: if $a_{p}(f)<0$, we could perform a twist by a character where $\chi(p)=-1$, instead. See Section 2.3 for a precise discussion.
@@ 1.6. Data analysis
Figure 3. 22-adic valuations of $\mathcal{L}$}-invariants in level $\Gamma_{0}(2)$ with respect to weights $10\leq k\leq 804$. (Total data: 13465 eigenforms.)
@@
modular of level $\Gamma_{0}(N)$, and its local restriction $\overline{r}=\overline{\rho}|_{\mathrm{Gal}(\overline{{{\mathbb{Q}}}}_{p}/{\mathbb{Q}}_{p})}$. For simplicity, assume $\overline{r}$ has only scalar endomorphisms. (See Section 6.2 for further discussion.) Given $\overline{r}$, a weight $k\geq 2$, and a sign $\pm$, we consider the set $X_{k}^{\pm}(\overline{r})\subseteq{\mathbb{P}}^{1}(\overline{{{\mathbb{Q}}}}_{p})$ consisting of $\mathcal{L}$}-invariants that arise from semi-stable, but non-crystalline, lifts of $\overline{r}$ with weight $k$ and sign $\pm$. The set $X_{k}^{\pm}(\overline{r})$ is the $\overline{{{\mathbb{Q}}}}_{p}$-points of the rigid generic fiber of a certain deformation ring, embedded into ${\mathbb{P}}^{1}$ by means of the $\mathcal{L}$}-invariant.
@@ 1.7. The distributional conjecture
$\mathbf{y}_{T}^{\pm}(\overline{\rho})=\left\{\frac{2(p+1)}{k(p-1)}\cdot v_{p}({\mathcal{L}}_{f})~\Big|~\parbox{135.15059pt}{\centering$f\in S_{k}(\Gamma_{0})^{\pm}$ is an eigenform, $k\leq T$, and $\overline{\rho}_{f}\simeq\overline{\rho}$\@add@centering}\right\}.$
Then, $\mathbf{y}_{T}^{\pm}(\overline{\rho})$ becomes equidistributed for Lebesgue measure on [1,0][-1,0] as $T\rightarrow\infty$.
When $p=2$, the normalization in Conjecture 1.3 is by $\frac{6}{k}$. (The slope in Figure 3 is about $\frac{1}{6}$.) Figure 4 presents a scatter plot and histogram to support Conjecture 1.3. In level $\operatorname{SL}_{2}({\mathbb{Z}})$, the only mod $p=2$ Galois representation is $\overline{\rho}=1\oplus 1$. We do not filter by the sign $\pm$ in the plot itself. See Section 8.3 for more plots.
@@ 1.10. Acknowledgments
The research reported on here was carried out, partially, while the authors visited the Max-Planck Institute for Mathematics in Bonn, Germany. The staff and members of the institute are warmly thanked for their support and hospitality. Bergdall’s research has been partially supported by Simons Foundation Travel Support Grant for Mathematicians MPS-TSM-00713782 and National Science Foundation grant DMS-2302284. Pollack’s research has been partially supported by National Science Foundation grants DMS-1702178 and DMS-2302285 and by Simons Foundation Travel Support Grant for Mathematicians MPS-TSM-00002405.
@@ 2.1. Notations
We fix an isomorphism $\iota:{\mathbb{C}}\simeq\overline{{{\mathbb{Q}}}}_{p}$, thus allowing us to view an eigenform as defined over $\overline{{{\mathbb{Q}}}}_{p}$. On $\overline{{{\mathbb{Q}}}}_{p}$, we also fix a $p$-adic valuation $v_{p}$ such that $v_{p}(p)=1$. We always measure $p$-adic numbers with $v_{p}(-)$. If we apply $v_{p}(-)$ to a complex number, we mean to implicitly apply $\iota$ first.
@@ 2.2. Modular symbols and $p$-adic $L$-functions
$\phi_{f}(P,r):=2\pi i\int_{\infty}^{r}f(z)P(z)~\!dz$
where $r\in{\mathbb{Q}}$ and $P(z)$ is an polynomial over $\mathbb{Z}$} of degree at most k2k-2 (as in [57]). The modular integrals of an eigenform $f$ may be decomposed into plus and minus parts and renormalized to take algebraic values:
@@
(3) $\int_{a+p^{n}{\mathbb{Z}}_{p}}P(x)~\!d\mu_{f}(x)=\mu_{f}(P,a,p^{n})$
for all $P(z)$ of degree less than or equal to k2k-2. The distribution $\mu_{f}$ is the $p$-adic $L$-function of $f$.
We note that $p$-adic $L$-functions can also be defined for eigenforms that are old at $p$, but in that case one also needs to choose a root of the Hecke polynomial $x^{2}-a_{p}(f)x+p^{k-1}$. When this root has valuation strictly less than k1k-1, then the construction of $\mu_{f}$ follows the same approach except that the definition of $\mu_{f}$ in (2) is slightly different and depends upon this choice of root. We will not discuss $p$-adic $L$-functions of forms that are old at $p$ any further.
Let us describe the $p$-adic $L$-function in the “$s$-variable”. First, write each $x\in{\mathbb{Z}}_{p}$ uniquely as $x=\langle x\rangle\omega(x)$, where $\langle x\rangle\in 1+p{\mathbb{Z}}_{p}$ and $\omega(x)$ is a (p1)(p-1)-st root of unity. Then, for a tame character $\psi:{\mathbb{Z}}_{p}^{\times}\to{\mathbb{F}}_{p}^{\times}\to\overline{{{\mathbb{Q}}}}_{p}^{\times}$, we set
$L_{p}(f,\psi,s):=\int_{{\mathbb{Z}}_{p}^{\times}}\psi(x)\langle x\rangle^{s-1}~\!d\mu_{f}(x).$
This is the $\psi$-branch” of the $p$-adic $L$-function and here $\psi$ has p1p-1 choices: $1$, $\omega$, $\omega^{2}$, $\dots$, $\omega^{p-2}$. Nonetheless, for $j$ an integer, we often abuse notation and simply write
$L_{p}(f,j)=\int_{{\mathbb{Z}}_{p}^{\times}}x^{j-1}~\!d\mu_{f}(x)$
@@
$\int_{{\mathbb{Z}}_{p}^{\times}}x^{j-1}~\!d\mu_{f}(x)=\int_{{\mathbb{Z}}_{p}^{\times}}\omega(x)^{j-1}\langle x\rangle^{j-1}~\!d\mu_{f}(x)=L_{p}(f,\omega^{j-1},j)$
@@
(4) $L_{p}^{\prime}(f,j):=\frac{d}{ds}\bigg|_{s=j}\left(\int_{{\mathbb{Z}}_{p}^{\times}}\omega(x)^{j-1}\langle x\rangle^{s-1}d\mu_{f}(x)\right).$
@@
$v_{p}\left(\int_{a+p^{n}{\mathbb{Z}}_{p}}(x-a)^{m}~\!d\mu_{f}(x)\right)\geq n\left(m-\frac{k-2}{2}\right).$
@@
This divisibility is almost [57, pg. 13, III], but in that reference a lower bound of $n\left(m-\frac{k-2}{2}\right)-\frac{k-2}{2}$ is given. However, in our situation of a form whose level is divisible by $p$ exactly once, we can achieve the slightly stronger bound of this lemma. Indeed, if $f$ had level prime-to-$p$ and $\alpha$ was a root of the Hecke polynomial of $f$ at $p$, then the very definition of $\mu_{f}$ yields that $\alpha^{n+1}\mu_{f}(P,a,p^{n})$ is integral. However, in our case, (2) implies that $\alpha^{n}\mu_{f}(P,a,p^{n})$ is integral where $\alpha=a_{p}(f)=\pm p^{\frac{k}{2}-1}$. Tracing through the proof of [57, pg. 13, III] using this stronger fact yields the claim of this lemma. ∎